What is the five-year discount factor given a 15% annual discount rate?

  • 0.497

  • 0.750

  • 0.625

  • 0.250
Answer #1

Given,

Annual discount rate = 15% or 0.15

Solution :-

Five-year discount factor = 1/(1 + annual discount factor)5

= 1/(1 + 0.15)5

= 1/(1.15)5

= 1/2.011357 = 0.497